OpenStudy (anonymous):

f(x)=8^(5sin^2(x+1)) then f'(pi/4)

OpenStudy (anonymous):

Its a derivative.

OpenStudy (anonymous):

actually i think you need the derivative take the log

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

\[\frac{d}{dx}b^x=b^x\ln(b)\] so by the chain rule \[\frac{d}{dx}b^{f(x)}=f'(x)b^{f(x)}\ln(b)\]